# How iteration composition influences convergence and stability in deep learning

### v2

@@ 2 A backward contraction principle
As for the condition in ˜2, let us verify that it is satisfied for SGD in the case of an overparametrized model like a neural network. Namely, in this setting, we generally consider losses of the form $L(\theta)=\frac{1}{N}\sum_{i=1}^{N}L_{i}(\theta)$, where $L_{i}$ is the loss computed on batch $B_{i}$ with $L_{i}(\theta)$ being non-negative. Because of the over-parameterization, there generally exists points $\theta^{*}$ where the full loss vanishes. This in turns implies that $L_{i}(\theta^{*})=0$ for all $i$. For SGD, the operators are $T_{i}(\theta)=\theta-h\nabla L_{i}(\theta)$. Now it is easy to see that ˜2 is satisfied: $d(\theta^{*},T_{i}(\theta^{*}))=\|\theta^{*}-\theta^{*}-h\nabla L_{i}(\theta^{*})\|=0,$ since the global zero $\theta^{*}$ is a critical point for all the $L_{i}$’s.
Now here is a major point: The contraction mapping principle fails with the forward order. As the next example demonstrates, it is not enough for the operator $T_{i}$’s to be uniform contractions for the forward sequence of iterates to converge, while this is always true for the backward sequence because of Theorem˜2.2.
@@
(Forward iterations counter-example.) Here is an extreme example illustrating the convergence failure for the forward sequence of iterates (even when the maps are uniform contractions), while the backward sequence converges to a single point under the conditions of Theorem˜2.2. Consider the two constant maps $S(\theta)=x_{0}$, and $U(\theta)=y_{0}$, which are contractions with $k=0$. Assume that $x_{0}\neq y_{0}$, now also assume that each $T_{i}$ is either equal to $S$ or $U$ with equal probability to be selected. Then, independently of $\theta$, the forward iterates
@@ 2.2 Relation between forward, backward, and generalization
First of all, recall that a map $T:\Omega\rightarrow\Omega$ produces a corresponding push-forward map $T_{*}:P(\Omega)\rightarrow P(\Omega)$ at the level of probability measures $P(\Omega)$ on $\Omega$. Recall that a probability measure $\mu\in P(\Omega)$ associates to each subset $A\subset\Omega$ a number $\mu(A)$ modelling the probability $\textrm{Prob}_{\mu}(\theta\in A)$ of $\theta$ being in the subset $A$ if $\theta$ is sampled from $\mu$. Now, if $\mu$ is a probability measure on $\Omega$ then the push forward $T_{*}\mu(A)=\mu(T^{-1}(A))$ for $A\subset\Omega$ computes the probability $\textrm{Prob}_{\mu}(T(\theta)\in A)$ of $T(\theta)$ being in $A$ if $\theta$ is sampled according to $\mu$.
Therefore, the forward sequence $(T_{n}\cdots T_{2}T_{1}):\Omega\rightarrow\Omega$ induces a push-forward $(T_{n}\cdots T_{2}T_{1})_{*}$ that maps an initial distribution $\mu_{0}$ to the probability measure $\mu_{n}=(T_{n}\cdots T_{2}T_{1})_{*}\mu_{0}$ modeling the probability distribution of the $n^{th}$ forward iterate $\theta_{n}=T_{n}\cdots T_{2}T_{1}(\theta_{0})$ when the initial point $\theta_{0}$ is randomly sampled from $\mu_{0}$. Observe that the forward iterates $\theta_{n}$ form a sequence of random variables, which is a Markov chain (since for the forward iterates we have that $\theta_{n}=T_{n}(\theta_{n-1})$). Many previous works (see related work in Section˜1.1) have shown under different assumptions that this sequence of forward iterates do not converge to a point when using fixed learning rates, but rather their probability distributions $\mu_{n}$ converge to a limiting probability distribution, called the stationary distribution of the Markov chain. In other words, starting at initial point $\theta_{0}$ the distribution $\mu_{n}=(T_{n})_{*}\cdots(T_{1})_{*}\delta_{\theta_{0}}$ of the forward iterate $\theta_{n}$ converges to a stationary distribution $\mu_{\theta^{*}}$. Recall that $\delta_{\theta_{0}}$ is the Dirac measure concentrated at $\theta_{0}$ (i.e., $\delta_{\theta_{0}}(A)$ is $1$ if $A$ contains $\theta_{0}$ and 0 otherwise).
@@ Appendix C Toward an approximate backward SGD
Consider the SGD operators $T_{i}(\theta)=\theta-h\nabla L_{i}(\theta)$ obtained by taking the gradient of a loss function on a batch $B_{i}$ of data at step $i$. We denote by $\theta_{n}=T_{n}\cdots T_{1}(\theta_{0})=\theta_{n-1}-h\nabla L_{n}(\theta_{n-1})$ the forward SGD iterate starting at initial point $\theta_{0}$ and by $\theta_{n}^{B}=T_{1}\cdots T_{n}(\theta_{0})$ the corresponding backward iterate starting at the same initial point. Motivated by Theorem˜C.2, we introduce the approximate backward iterate $\tilde{\theta}_{n}$ as follows:
@@ D.2.1 Experiments
We trained a MLP with 5 layers of 500 neurons each with stochastic gradient descent with no regularization to learn the 10 classes of Fashion-MNIST dataset Xiao et al. (2017). We used a learning rate of $h=0.001$ (and $h^{\prime}=h/c$), and a batch-size of |B|=2048|B|=2048222Or the closest integer divisible by $c$; e.g., $2049$ for $c=3$.. We examine three different shrinking factors $c\in\{2,3,4\}$. In each setting, we test the values of $\lambda\in\{0.5h^{2},h^{2},2h^{2},4h^{2}\}$. The following figures present the train and test loss over training compared to training with large-batch $|B|=2048$ and learning rate $h^{\prime}=h/c$, and training with small-batch $2048/c$ with learning rate $h^{\prime}$. The experiment shows that our method can outperform vanilla training both for large and small batches with minimal tuning of $\lambda$ for this specific experiment.

### v3

@@ 2 A backward contraction principle
As for the condition in ˜2, let us verify that it is satisfied for SGD in the case of an overparametrized model like a neural network. Namely, in this setting, we generally consider losses of the form $L(\theta)=\frac{1}{N}\sum_{i=1}^{N}L_{i}(\theta)$, where $L_{i}$ is the loss computed on batch $B_{i}$ with $L_{i}(\theta)$ being non-negative. Because of the over-parameterization, there generally exists points $\theta^{*}$ where the full loss vanishes. This in turns implies that $L_{i}(\theta^{*})=0$ for all ii. For SGD, the operators are $T_{i}(\theta)=\theta-h\nabla L_{i}(\theta)$. Now it is easy to see that ˜2 is satisfied: $d(\theta^{*},T_{i}(\theta^{*}))=\|\theta^{*}-\theta^{*}-h\nabla L_{i}(\theta^{*})\|=0,$ since the global zero $\theta^{*}$ is a critical point for all the $L_{i}$’s.
Now here is a major point: The contraction mapping principle fails with the forward order. As the next example demonstrates, it is not enough for the operator $T_{i}$’s to be uniform contractions for the forward sequence of iterates to converge, while this is always true for the backward sequence because of Theorem˜$2.$
@@
(Forward iterations counter-example.) Here is an extreme example illustrating the convergence failure for the forward sequence of iterates (even when the maps are uniform contractions), while the backward sequence converges to a single point under the conditions of Theorem˜$2.$ Consider the two constant maps $S(\theta)=x_{0}$, and $U(\theta)=y_{0}$, which are contractions with $k=0$. Assume that $x_{0}\neq y_{0}$, now also assume that each $T_{i}$ is either equal to $S$ or $U$ with equal probability to be selected. Then, independently of $\theta$, the forward iterates
@@ 2.2 Relation between forward, backward, and generalization
First of all, recall that a map $T:\Omega\rightarrow\Omega$ produces a corresponding push-forward map $T_{*}:P(\Omega)\rightarrow P(\Omega)$ at the level of probability measures $P(\Omega)$ on $\Omega$. Recall that a probability measure $\mu\in P(\Omega)$ associates to each subset $A\subset\Omega$ a number $\mu(A)$ modelling the probability $\textrm{Prob}_{\mu}(\theta\in A)$ of $\theta$ being in the subset AA if $\theta$ is sampled from $\mu$. Now, if $\mu$ is a probability measure on $\Omega$ then the push forward $T_{*}\mu(A)=\mu(T^{-1}(A))$ for $A\subset\Omega$ computes the probability $\textrm{Prob}_{\mu}(T(\theta)\in A)$ of $T(\theta)$ being in AA if $\theta$ is sampled according to $\mu$.
Therefore, the forward sequence $(T_{n}\cdots T_{2}T_{1}):\Omega\rightarrow\Omega$ induces a push-forward $(T_{n}\cdots T_{2}T_{1})_{*}$ that maps an initial distribution $\mu_{0}$ to the probability measure $\mu_{n}=(T_{n}\cdots T_{2}T_{1})_{*}\mu_{0}$ modeling the probability distribution of the $n^{th}$ forward iterate $\theta_{n}=T_{n}\cdots T_{2}T_{1}(\theta_{0})$ when the initial point $\theta_{0}$ is randomly sampled from $\mu_{0}$. Observe that the forward iterates $\theta_{n}$ form a sequence of random variables, which is a Markov chain (since for the forward iterates we have that $\theta_{n}=T_{n}(\theta_{n-1})$). Many previous works (see related work in Section˜1.1) have shown under different assumptions that this sequence of forward iterates do not converge to a point when using fixed learning rates, but rather their probability distributions $\mu_{n}$ converge to a limiting probability distribution, called the stationary distribution of the Markov chain. In other words, starting at initial point $\theta_{0}$ the distribution $\mu_{n}=(T_{n})_{*}\cdots(T_{1})_{*}\delta_{\theta_{0}}$ of the forward iterate $\theta_{n}$ converges to a stationary distribution $\mu_{\theta^{*}}$. Recall that $\delta_{\theta_{0}}$ is the Dirac measure concentrated at $\theta_{0}$ (i.e., $\delta_{\theta_{0}}(A)$ is 11 if AA contains $\theta_{0}$ and 0 otherwise).
@@ Appendix C Toward an approximate backward SGD
Consider the SGD operators $T_{i}(\theta)=\theta-h\nabla L_{i}(\theta)$ obtained by taking the gradient of a loss function on a batch $B_{i}$ of data at step ii. We denote by $\theta_{n}=T_{n}\cdots T_{1}(\theta_{0})=\theta_{n-1}-h\nabla L_{n}(\theta_{n-1})$ the forward SGD iterate starting at initial point $\theta_{0}$ and by $\theta_{n}^{B}=T_{1}\cdots T_{n}(\theta_{0})$ the corresponding backward iterate starting at the same initial point. Motivated by Theorem˜C.2, we introduce the approximate backward iterate $\tilde{\theta}_{n}$ as follows:
@@ D.2.1 Experiments
We trained a MLP with 5 layers of 500 neurons each with stochastic gradient descent with no regularization to learn the 10 classes of Fashion-MNIST dataset Xiao et al. (2017). We used a learning rate of $h=0.001$ (and $h^{\prime}=h/c$), and a batch-size of |B|=2048|B|=2048222Or the closest integer divisible by $c$; e.g., 20492049 for $c=3$.. We examine three different shrinking factors $c\in\{2,3,4\}$. In each setting, we test the values of $\lambda\in\{0.5h^{2},h^{2},2h^{2},4h^{2}\}$. The following figures present the train and test loss over training compared to training with large-batch $|B|=2048$ and learning rate $h^{\prime}=h/c$, and training with small-batch $2048/c$ with learning rate $h^{\prime}$. The experiment shows that our method can outperform vanilla training both for large and small batches with minimal tuning of $\lambda$ for this specific experiment.